Find the geometric means of each sequence
1. 9,___, ____, -243
2. 1.5, ___, ___, ___ 384

Answer by stanbon(75887)

Find the geometric means of each sequence
1. 9,___, ____, -243
a=9
ar^3=-243
9r^3=-243
r^3=-27
r=-3
a(2)=9*-3=-27
a(3)=-37*-3=81
The sequence: 9,-27,81,-243
----------------------
2. 1.5, ___, ___, ___ 384
a=1.5
a(5)=ar^4
ar^4=1.5r^4=384
r^4=884/1.5=256
r=4
Sequence: 1.5, 1.5*4=6,6*4=24,24*4=96,96*4=384
Cheers,
Stan H.